OpenStudy (anonymous):

quadrilateral ABCD ~ quadrilateral WXYZ, AB= 15, BC= 27, BC=27 and the scale factor of WXYZ to ABCD is 2/3. find XY. HELP!

OpenStudy (anonymous):

Corresponding parts of your quadrilateral are proportional. So AB/WX = BC/XY = CD/YZ = DA/ZW. And they are all equal to 2/3, the scale factor. So BC/XY = 2/3. So find out XY from there since you know BC.